The following information relates to an exchange of assets by Wharton Company. The exchange lacks commercial substance.

Old Equipment

Book Value

Fair Value

Cash Paid

Case I

$75,000

$85,000

$15,000

Case II

$50,000

$45,000

$7,000

For Case II, Wharton records the equipment at $ Answer on its books and gain or (loss) of $ Answer. NB. if reporting a loss, use minus sign in answer.

On January 2, 2015, Deck Inc. purchased machinery with a cost of $10,440,000, a useful life of 10 years and no salvage value. The company uses straight-line depreciation. At December 31, 2015 and December 31, 2016, the company determines that impairment indicators are present. The following information is available for impairment testing at each year end:

  12/31/2015                  12/31/2016

            Fair value less costs to sell                  $9,315,000                  $8,850,000

            Value-in-use                                        $9,350,000                  $8,915,000

           

There is no change in the asset’s useful life or salvage value. The 2016 income statement will report (impairment loss or recovery of impairment loss) Answer of $ Answer

January, 2016, Gems Imperial purchased a mineral mine for $3,400,000 with removable ore estimated by geological surveys at 2,000,000 tons. The property has an estimated value of $200,000 after the ore has been extracted. The company incurred $1,000,000 of development costs preparing the mine for production. During 2016, 500,000 tons were removed and 400,000 tons were sold. What is the amount of depletion that Gems Imperial should expense for 2016? $
Answer #1

The answer to the first question will be as follows,

In quention it is given that, the exchange lacks commercial substance. Then the entry will be as follows

New Equipment Dr 45000

Loss on Exchange Dr 12000

To Cash 7000

To Old Equipment 50000

Thus, For Case II, Wharton records the Equipement at $45000 in its books and loss of ($12000) or -$12000.

Answer to the Second question:

Given facts in the question are as follows:

Cost of Machinery = $10440000

Useful life = 10 years

Salvage value = NIL

Hence Depreciation for 2015 = (Cost - Salvage value)/ Useful life

= ($10440000-0)/10 = $1044000 per annum

The carrying amount of the machine as on Dec 31, 2015 = Cost - Depreciation

= 10440000-1044000 = $9396000

The Fair value of the machine net of cost to sell as on Dec 31, 2015 .ie, Net Realisable value = $9315000

The value in use as on dec 31, 2015 =$ 9350000

The Recoverable amount is higher of NRV or Value in use = $9350000

Here. recoverable amount of the machine is $9350000 but the carrying amount of the machine in balance sheet is $9396000 which is not fair.

Hence, the entity requires to bring the asset value to its recoverable amount by impairing $46000

The Journal entry for the impairement loss is

Impairment loss account (To be charged to Profit and loss account) Dr $ 46000

To Machine account $46000

Then,

Balance sheet as on Dec 31, 2015:

Machinery 9396000

Less : Impairment Loss 46000

Carrying amount 9350000

This carrying amount of 9350000 should be depreciated over the remaining useful life of the machine .ie, over 9 years.

Then, Depreciation for 2016 = 9350000/9 years = 1038888 per annum

Carrying amount as on Dec 31, 2016 = 9350000-1038888 = 8311112

NRV as on Dec 31, 2016 = 8850000

Value in use as on Dec 31, 2016 = 8915000

Recoverable amount = higher of NRV and value in use = 8915000

As carring amount is less and Recoverable amount is more, we have to reverse the Impairment loss recignised in earlier year. Then the entry will be

Machine account Dr 46000

To Impairment loss account (To be transfered to Profit and loss account) 46000

Note: The Reversal sholu be restricted to the Impairment loss trasfered to profit and loss account in earlier years.

Hence, The 2016 Income statement will report recovery of impairement loss of $ 46000

The answer for third question is :

The Cost of the Mine = $3400000

Add: Developement Cost = $1000000

Total Cost of Mine =$ 4400000

The Salvage value of Mine = $200000

The total Estimated Minerals capacity of Mine = 2000000 Tons

The Extracted tons during 2016 = 500000

Depletion Rate = (Cost - Salvage value)/ Estimated capacity

= (4400000 - 200000)/2000000 =$ 2.10 per ton

Total Depletion of Mine for 2016 = Rate per ton * Extacted tons = 2.10* 500000 = $1050000

But , This total depletion shall not be recorded as Depletion Expense. It is bacauese, the Gem Imperal sold only 400000 tons of minerals and there is still 100000 tons of minerals are in stock at the end of 2016.

Thus, Depletion Expense to be recognised in 2016 = 2.10 * 400000 = $840000

The Entry at the end of 2016 will be

Mineral Inventory Dr $210000

Depletion Expense Dr $ 840000

To Mine asset $1050000